OpenStudy (anonymous):

let f be the function given by f(x)=cos(x). if the number c satisfies the conclustion of the mean value theorem for f on the clsoed interval [2,4], then c=

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

you have to solve \[-\sin(c)=\frac{\cos(4)-\cos(2)}{4-2}\] but other than 4 - 2 = 2 i would have no idea how to do it without technology

OpenStudy (anonymous):

did that, calculators can be used. I used my calculator to graph y=sin(x) and then y=.11875 and got 3.023 and another person did the same thing with the same rounding and everything and got 3.142, but both numbers are options

OpenStudy (anonymous):

oh and i used the intersection tool on the calculator to get that answer